OpenStudy (anonymous):

Bats use ultrasound in echolocation. This sound is characterized by _____. low frequency low amplitude high frequency high amplitude

OpenStudy (anonymous):

A galloping pony speeds past you at 5 m/s. The frequency of the sound produced by the hooves on the dirt is 221 Hz. Assume the speed of sound to be 342 m/s. What is the observed frequency of the hooves after the pony has passed your position? 218 Hz 226 Hz 216 Hz 224 Hz

OpenStudy (sidsiddhartha):

high frequency

OpenStudy (sidsiddhartha):

frequency at receiver F.R. is given as, F.R. = fs*C/(C+Vs) = 221*342/(342+5)=218

OpenStudy (anonymous):

thanks :)

OpenStudy (sidsiddhartha):

:)